2.(2 points) The police department writes parking tickets (at random, regardless of guilt) to 20% of all cars parked on any faculty/staff parking lot on any given day.

a) What is the average/mean number of tickets written on any given day in a parking lot with 13 cars.

ANSWER: Average = _______________________________________ (simplified number)

b) Find the probability that exactly 4 tickets will be written tomorrow on lot A and on lot B, each of which fits 13 cars. State your assumptions

ANSWER: Pb = _______________________________________ (simplified number)
Answer #1

Given that the police department writes parking tickets.

P=20% =0.2

(a) n=13

average/ mean nukmber of tickets written = np = 13 \times 0.2 = 2.6

(b) n= 13 + 13 = 26

P=0.2

P(x=4) = 26c4 (0.2)4 (0.8)22 = 0.1765